OpenStudy (anonymous):

Find the exact value using sum and difference formulas: tan75

OpenStudy (campbell_st):

try tan(75) = tan(30 + 45) and using the general for for the sum in tan its \[\tan(A +B) = \frac{\tan(A) + \tan(B)}{1 - \tan(A)\tan(B)}\] both angles have exact values... hope this helps.

OpenStudy (anonymous):

I keep getting stuck after I plug everything in and write their real values

OpenStudy (campbell_st):

ok so \[\tan(30) = \frac{1}{\sqrt{3}}...and....\tan(45) = 1\] so you get \[\frac{\frac{1}{\sqrt{3}} + 1}{1 - \frac{1}{\sqrt{3}} \times 1}\] so it looks like you have \[\frac{1 + \sqrt{3}}{\sqrt{3}} \div \frac{\sqrt{3} -1}{\sqrt{3}}\] flip and multiply since you are dividing by a fraction gives \[\frac{1 + \sqrt{3}}{\sqrt{3} -1}\] I'd say you would need to rationalize the denominator to get the final answer.

OpenStudy (anonymous):

thank you :)

OpenStudy (campbell_st):

glad to help